OpenStudy (anonymous):

limit question as x approaches sqrt(x^2+1) find the limit

OpenStudy (anonymous):

as x approaches 3? 0? \(\infty)?

OpenStudy (anonymous):

as x approaches infinity

OpenStudy (anonymous):

sqrt(x^2+1)

OpenStudy (anonymous):

as x gets larger, so does \(\sqrt{x^2+1}\)

OpenStudy (anonymous):

so \(\infty\)

OpenStudy (anonymous):

so the answer is just infinity?

OpenStudy (anonymous):

of course \(x\) gets bigger, \(x^2+1\) gets bigger, \(\sqrt{x^2+1}\) gets bigger think of what the graph looks like after a while it just looks like the line \(y=x\)

OpenStudy (anonymous):

hmmm thanks